OpenStudy (anonymous):

Solve. 8/x + 15/x^2 = -1

OpenStudy (anonymous):



OpenStudy (anonymous):

x^2+5x+15=0 can u solve it ?

OpenStudy (anonymous):

no i can't

OpenStudy (anonymous):

ok i'll help u .

OpenStudy (anonymous):

first i multiple the equation by x^2 ok ?

OpenStudy (anonymous):

wht im confused can you do it step by step for me

OpenStudy (anonymous):

then when we have a polynomial like ax^2+bx+c =0 we can find the roots with this method sqrt(b^2-4ac) = sqrt ( delta ) roots = (-b+-sqrt ( delta ) )/2a

OpenStudy (anonymous):

let me scan the answer .

OpenStudy (anonymous):

wait 2 min

OpenStudy (anonymous):

did u get it ?

OpenStudy (anonymous):

yes thankss

OpenStudy (anonymous):

u'r welcom

OpenStudy (anonymous):

wait im sorry your hand writing is kind of confusing .. do u think u cn write the whole thing over